If the top of a coke can has a 7 cm diameter, what is the radius, circumfrence and area?

Answer by rfer(16322)

r=.5d
r=3.5 cm
-----------
c=pi*d
c=7*3.1416
c=22 cm
--------------
a=pi*r^2
a=3.14*3.5^2
a=3.14*12.25
a=38.47 sq cm